# Progressive Quantum Algorithm for Maximum Independent Set with Quantum Alternating Operator Ansatz

### v2

@@ IV.1 The effects of graph construction rules and $n_{0}$ on the performance of PQA
(c) OAR, 2-regular graphs
(d) OAR, 3-regular graphs
@@
Figure 8: The AR achieved by PQA, QLS, and DS-QAOA+ under the same level depth and optimization runs. In each subplot, the curves depict the AR achieved by the different algorithms across various graph types as the level depth $p$ increases.
IV.2 The performance comparison among different algorithms
IV.2.1 The OAR obtained by different algorithms under the same number of optimization runs and level depth.
In our simulations, we randomly generate 20 graphs each for different ER graphs (with probability = 0.4, 0.5) and $k$-regular graphs ($k=2,3$) with the number of nodes $n=14$. Subsequently, we respectively test the performance of PQA, QLS, and DS-QAOA+ on these graphs. For PQA, the tolerance on function values $\xi=0.1$ and the quantum state $|s\rangle$ is initialized as $|\mathbb{0}\rangle$ at each solving process. The classical outer parameter optimization loop ends when the convergence tolerance Ni et al. (2024) satisfies $\varepsilon\leq 0.001$. We set the level depth from one to five, and each algorithm is executed $100\times p$ optimization runs at each depth. At the end of the runs, we calculate the OAR achieved by each algorithm. These experimental procedures yield 20 sets of OAR for each graph type. The mean value of these sets provides the average OAR for various algorithms at depth $p$. From these numerical results illustrated in FIG. 8, we observe that the achieved OAR tends to increase as the level depth increases, thereby enhancing the quality of the obtained solution. More importantly, these numerical results demonstrate that the OAR achieved by PQA is superior to that obtained by DS-QAOA+ and QLS at the same level depth and number of optimization runs. More specifically, on the ER graphs with prob = 0.4, the OAR achieved by PQA is approximately $8.24\%$ ($13.668\%$) higher than that achieved by QLS (DS-QAOA+). On the ER graphs with prob = 0.5, the OAR achieved by PQA is approximately $8.128\%$ ($14.422\%$) higher than that achieved by QLS (DS-QAOA+). On the 2-regular graphs, the OAR achieved by PQA is about $3.612\%$ ($4.514\%$) higher than that achieved by QLS (DS-QAOA+). On the 3-regular graphs, the OAR obtained by PQA is about $8.486\%$ and $9.438\%$ higher than that achieved by QLS and DS-QAOA+, respectively.
Table 3: The required level depths (RLDs), the total number of optimization runs (TRs) and iterations (TITRs), total qubits (T-qubits), total circuit depths (TCDs), total multi-qubit CR (T-MQCR) gates, and total quantum (T-Q) gates consumed by different algorithms on specific graph types when they achieve OAR of 0.95.
Types Algo. RLDs TRs TITRs T-qubits TCDs T-MQCR gates T-Q gates
ER, prob = 0.4 DS-QAOA+ 4.9 91.6192 5664.43 1282.669 19304.17 6285.077 12570.15
QLS 4.4 26.8412 1666 371.46 5282.35 1648.8952 3297.79
PQA 1.35 17.7572 1652 112.76 464.35 135.84 337.3868
ER, prob = 0.5 DS-QAOA+ 6.85 339.1095 21624 4747.533 99884.7 32520.6 65041.2
QLS 4.55 34.3853 1939.38 478.59 7011.16 2186.785 4373.57
PQA 1.6 15.8492 1241.81 110.94 527.78 160.869 361.362
2-regular DS-QAOA+ 3.35 10.4152 735.41 145.813 1500.31 488.473 976.946
QLS 2.35 5.35 374.34 50.02 655.375 202.498 405
PQA 1 5.15 336.96 41.74 91.89 22.497 83.48
3-regular DS-QAOA+ 6.3 493.9459 36313 6915.243 133809.9 43566.03 87132.06
QLS 6.21 50.3033 3519.34 690.37 14005.5 4359.867 8719.735
PQA 1.35 20.5773 2319.783 150.21 482.54 125.522 407.431
Table 4: The ratio of quantum resources saved by PQA relative to DS-QAOA+ and QLS when an OAR of 0.95 is achieved.
Types Algo. RLDs TRs TITRs T-qubits TCDs T-MQCR gates T-Q gates
ER, prob = 0.4 DS-QAOA+ 0.7245 0.8062 0.7083 0.9121 0.9759 0.9784 0.9732
QLS 0.6932 0.3384 0.0084 0.6964 0.9121 0.9176 0.8977
ER, prob = 0.5 DS-QAOA+ 0.7664 0.9533 0.943 0.9766 0.9947 0.9951 0.9944
QLS 0.6484 0.5391 0.36 0.7682 0.9247 0.9264 0.9174
2-regular DS-QAOA+ 0.7015 0.5054 0.5418 0.7137 0.9388 0.9539 0.9145
QLS 0.5745 0.0379 0.1 0.1655 0.8598 0.8889 0.7939
3-regular DS-QAOA+ 0.7857 0.9583 0.9361 0.9783 0.9964 0.9971 0.9953
QLS 0.7826 0.5909 0.3408 0.7824 0.9655 0.9712 0.9533
@@ IV.2 The performance comparison among different algorithms
IV.2.2 The quantum resources consumed by various algorithms when they achieve the same OAR.
To provide a more comprehensive comparison of the performance of DS-QAOA+, QLS, and PQA, we further investigate the total number of optimization runs and iterations required by each algorithm to reach a given OAR at the required level depths. Moreover, for each graph instance, the total number of qubits, total circuit depths, and other resources demanded by each algorithm are also calculated. These indicators are used to clearly demonstrate the resource consumption of each algorithm when they reach the same OAR. In the numerical simulation experiments, we set the level depth from one to $n$, and the OAR obtained by each algorithm under multiple runs to be at least $0.95$. We perform 100 optimization runs at each level depth $p$ when $p\leq 5$, and $2^{p}$ optimization runs when $p>5$. For each graph instance, we calculate the RLDs, TRs, ITRs, and other resources when they achieve the OAR of 0.95. Averaging the data of the same graph type provides an estimate of the resources consumed by different algorithms when solving the MIS problem on various graph types to achieve the given OAR. The specific numerical results are shown in Table 3. These results suggest that PQA is usually able to achieve the OAR of 0.95 with fewer resources. Table 4 summarizes the quantum resources saved by PQA relative to DS-QAOA+ and QLS when an OAR of 0.95 is reached on a given graph type. For example, when solving the MIS problem on a 2-regular graph, the total qubits consumed by PQA is only $28.63\%$ of DS-QAOA+, saving $71.37\%$ in qubit usage. These results demonstrate that PQA not only achieves the given OAR of 0.95 in fewer optimization runs and iterations but also significantly reduces the consumption of qubits, circuit depth, and quantum gates compared with DS-QAOA+ and QLS, highlighting its efficiency. By minimizing the quantum resources required while maintaining high performance, PQA may be increasingly promising for practical implementation in scenarios where resource availability is limited. In our work, we also test other OAR thresholds (i.e., 0.8, 0.85, and 0.9) and show the corresponding resources required by each algorithm in the Appendix C.
(a) given OAR = 0.8
(b) given OAR = 0.85
(c) given OAR = 0.9
Figure 9: The required running time of DS-QAOA+, QLS, and PQA when they achieve the same given OAR.
IV.2.3 The comparison of the running time when DS-QAOA+, QLS, and PQA achieve the same OAR.
@@
To assess the performance of different algorithms in achieving the given OAR, we examine the resource consumption of QLS, PQA, and DS-QAOA+. Experimental results show that QLS and PQA typically require fewer resources than DS-QAOA+ to achieve the same OAR. For QLS and PQA, classical computations (i.e., determining whether an edge exists between two points) are involved in updating the global solution and graph expansion. In contrast, DS-QAOA+ excludes such computations. Although numerical simulations show that the time for a single edge judgment is minimal, we include this classical computation time in our runtime comparison for fairness. For QLS and PQA, the time required for a single optimization run includes both quantum and classical components. The classical time $T_{c}$ is the time consumed in determining whether there is an edge between vertices during updating the global solution or expanding the subgraph. Assuming that the time required for a single determination is $t_{j}$, and the total number of determinations required by the algorithm in one run is $TJ$, then
$T_{c}=TJ\times t_{j}.$ (14)
@@
The quantum running time $T_{q}$ includes the time required for initial state preparation, quantum state measurements, and quantum gate evolution across multiple iterations Guerreschi and Matsuura (2019). In each iteration, $M$ repeated measurements are needed to calculate the average expectation value of the objective function. The time consumed in a single repetition is
$t_{q}=t_{p}+t_{M}+(\text{circuit depth})\times t_{G},$ (15)
@@
$T_{q}=T_{ITR}\times\text{ITRs},$ (17)
where ITRs is the number of iterations required for the objective function to converge in this run. Therefore, the time consumed in a single run for QLS and PQA is
$T_{PQA}=T_{q,PQA}+T_{c,PQA},$ (18)
$T_{QLS}=T_{q,QLS}+T_{c,QLS},$ (19)

### v3

@@ IV.1 The AR obtained by various algorithms with the same number of optimization runs and level depths
(c) OAR, 3-regular graphs
(d) AAR, ER graphs with prob = 0.4
(e) AAR, ER graphs with prob = 0.5
(f) AAR, 3-regular graphs
@@
Figure 5: The AR achieved by various algorithms. The gray dotted line corresponds to the AR obtained by the hill climbing algorithm, and the purple dotted line corresponds to the AR obtained by the classical counterpart of PQA. In each subplot, the curves depict the AR achieved by different algorithms across various graph types with increasing level depth $p$.
The OAR is defined as the maximum approximation ratio obtained over multiple optimization runs. It represents the best performance of the algorithm and provides a measure of the highest-quality solution achieved in the experiments. The AAR is the average of the approximation ratios obtained in multiple optimization runs, providing a measure of stability. From these numerical results illustrated in Figure 5, we observe a clear trend that the achieved OAR increases as the level depth grows, indicating that the solutions obtained are closer to the global optimal solution with reduced approximation error. This improvement is primarily due to the increased ansatz depth, which enhances the expressive power of the quantum circuit, allowing it to represent a broader range of quantum states and enable a more effective search for high-quality solutions within the Hilbert space. Notably, the results also show that, given the same number of runs and level depths, PQA achieves a higher OAR and a significantly higher AAR compared with other quantum algorithms. This suggests that PQA not only excels in finding high-quality solutions but also maintains a more stable performance across multiple optimization runs. The higher AAR further implies that the solutions of PQA exhibit lower variance, ensuring that it reliably produces quasi-optimal solutions regardless of random parameter initialization. Such stability is crucial for practical applications as it reduces the risk of performance fluctuations, making PQA a more robust choice. The specific improvements are as follows.
  • On ER graphs with an edge probability of 0.4, the OAR achieved by PQA is approximately $8.240\%$ and $13.668\%$ higher than that achieved by QLS and DS-QAOA+, respectively. The AAR achieved by PQA is approximately $3.766\%$ and $4.416\%$ higher than that achieved by QLS and DS-QAOA+, respectively.
  • On ER graphs with an edge probability of 0.5, the OAR achieved by PQA is approximately $8.128\%$ and $14.422\%$ higher than that achieved by QLS and DS-QAOA+, respectively. The AAR achieved by PQA is approximately $6.048\%$ and $6.658\%$ higher than that achieved by QLS and DS-QAOA+, respectively.
  • On 3-regular graphs, the OAR obtained by PQA is approximately $8.486\%$ and $9.438\%$ higher than that achieved by QLS and DS-QAOA+, respectively. The AAR obtained by PQA is approximately $2.588\%$ and $1.618\%$ higher than that achieved by QLS and DS-QAOA+, respectively.
Compared with directly applying the hill-climbing algorithm to solve the original problem, the classical counterpart to PQA significantly improves the solution stability by progressively refining the obtained solution on each subgraph, demonstrating the effectiveness of our method in addressing the problem in an incremental manner. The results also show that the AR obtained by PQA can be higher than the classical counterpart to PQA, which benefits from quantum superposition and quantum parallelism. Specifically, the classical local search algorithm typically requires multiple verification and correction steps to guarantee solution feasibility QLS , and the feasible solution space constructed by the classical local search algorithm is inherently limited due to the sequential and local nature of its search process, as it evaluates and updates only a single solution per iteration. This limitation is mitigated in PQA, as it constructs a larger feasible solution space through quantum superposition and elaborate circuit designs. By leveraging quantum parallelism, QAOA+ evaluates multiple feasible solutions in parallel. Therefore, in PQA, QAOA+ is preferred over classical heuristic algorithms for solving the MIS problem on each subgraph.
@@
IV.2 The quantum resource consumption of various algorithms when achieving the same AR
In the previous subsection, we compared the obtained OAR and AAR of different algorithms given the same number of runs and level depths. However, comparing algorithms solely based on the same number of runs has limitations, as some algorithms may achieve a similar OAR with fewer runs, which could lead to an unfair assessment of their efficiency. To provide a more balanced evaluation, we further compare the total resource consumption of each algorithm when achieving the same given AR. We selected several key metrics to evaluate performance, including the number of iterations, the number of qubits, the number of multi-qubit controlled $R_{x}$ gates, and the algorithm runtime. By combining these metrics, we can more clearly and comprehensively reflect the resource consumption of different algorithms, offering valuable insights for performance optimization in practical applications. In Appendix B, we give the methods of calculating the total resources for each algorithm.
(a) Total iterations, ER with prob = 0.4
(b) Total iterations, ER with prob = 0.5
(c) Total iterations, 3-regular graphs
(d) Total runtime, ER with prob = 0.4
(e) Total runtime, ER with prob = 0.5
(f) Total runtime, 3-regular graphs
(g) Total qubits, ER with prob = 0.4
(h) Total qubits, ER with prob = 0.5
(i) Total qubits, 3-regular graphs
(j) Total multi-qubit gates, ER with prob = 0.4
(k) Total multi-qubit gates, ER with prob = 0.5
(l) Total multi-qubit gates, 3-regular graphs
Figure 6: The total resource consumption of various algorithms on different graph types.
FIG. 6 presents the total resource consumption required by different algorithms to achieve a given AR across various graph types. From the numerical results, we observe that the total resource consumption increases as the target AR increases, which aligns with expectations. This is because a higher AR requires the output solutions to be closer to the true solution, necessitating more optimization runs or deeper level depths to search for a better solution, increasing the overall resource consumption. However, DS-QAOA+ and PQA exhibit counterintuitive behavior in the resource consumption trends in some cases. Especially for DS-QAOA+, the total number of qubits and multi-qubit controlled $R_{x}$ gates for achieving AR = 0.9 is significantly higher than the corresponding consumption for achieving AR of 0.95. Upon examining the experimental data, we found that DS-QAOA+ reaches the target AR using fewer level depths but consumes more runs at an AR of 0.9. In contrast, at an AR of 0.95, DS-QAOA+ increases the number of level depths to achieve the given AR but reduces the number of runs required, as shown in FIG. A2. This phenomenon is related to the expressive ability of the PQC, which impacts the ability of the algorithm to find quasi-optima. Specifically, shallow level depths have limited expressive power and can explore only a smaller solution space during each optimization process. Therefore, to approach the target solution, the algorithm requires more runs to compensate for the limited search space explored. In contrast, deep level depths, with stronger expressive capabilities, can explore a broader solution space. This enables the algorithm to more precisely approach the target solution in each run, thus achieving the same AR with fewer runs. However, deeper level depths are not always better, the greater the level depth means the greater the demand for quantum gate and circuit depth, which may limit the scale of problem-solving on the current NISQ device. In our calculations, the total consumption is determined by the product of the total number of runs and the average resource consumption per run. While increasing level depth inevitably raises the number of gates per run, the increase in gate count is relatively small compared with the difference in the number of optimization runs between AR of 0.9 and AR of 0.95. Moreover, increasing the level depth does not lead to an increase in the number of qubits consumed per run. As a result, for AR = 0.9, the total number of multi-qubit controlled $R_{x}$ gates and qubits is higher compared with AR of 0.95. The above analysis can also explain some counterintuitive phenomena in PQA. From the numerical simulation results in FIG. 6, we also observed that for smaller target AR (e.g., 0.8 and 0.85), DS-QAOA+ and QLS typically achieve the given AR with fewer iterations and shorter runtime compared with PQA. As shown in FIG. A2, for AR values of 0.8 and 0.85, the total number of optimization runs consumed by DS-QAOA+, QLS, and PQA is similar. However, since PQA performs more optimization rounds per run than DS-QAOA+ and QLS, it results in a higher total iteration consumption. Moreover, the consumption of iterations has a significant impact on the overall runtime as shown in Appendix B.1, with more iterations generally leading to higher runtime consumption. Nevertheless, it is important to note that PQA consumes significantly fewer qubits and multi-qubit controlled gates compared with DS-QAOA+ and QLS. While the number of iterations is important, the number of qubits and gates required per run more directly determines the problem size that can be supported by resource-constrained quantum devices. Excessive qubit and gate demands per iteration may limit the ability of hardware to solve larger-scale problems efficiently. Therefore, sacrificing some iterations to reduce qubit and gate consumption is a justified trade-off, as it allows the algorithm to tackle larger problems within the constraints of current quantum hardware. However, this does not imply that PQA always requires the most iterations for any AR. Numerical results show that as the target AR increases, the efficiency of PQA becomes more evident. This is mainly because PQA can achieve the desired AR with fewer runs than DS-QAOA+ and QLS, significantly reducing overall resource consumption. By minimizing the quantum resources required while achieving high performance, PQA may be increasingly promising. Detailedly, for each graph type, the resource consumption of PQA compared with DS-QAOA+ and QLS for a given AR of 0.95 is as follows.
  • On the ER graphs with an edge probability of 0.4, the number of iterations consumed by PQA is approximately $29.17\%$ and $99.16\%$ of those consumed by DS-QAOA+ and QLS, respectively. The amount of runtime consumed by PQA is approximately $29.38\%$ and $93.48\%$ of those consumed by DS-QAOA+ and QLS, respectively. The number of qubits consumed by PQA is approximately $8.79\%$ and $30.36\%$ of those consumed by DS-QAOA+ and QLS, respectively. The number of multi-qubit controlled gates consumed is $2.16\%$ and $8.24\%$ of those consumed by DS-QAOA+ and QLS, respectively.
  • On the ER graphs with an edge probability of 0.5, the number of iterations consumed by PQA is approximately $5.70\%$ and $64.00\%$ of those consumed by DS-QAOA+ and QLS, respectively. The amount of runtime consumed by PQA is approximately $5.80\%$ and $59.95\%$ of those consumed by DS-QAOA+ and QLS, respectively. The number of qubits consumed by PQA is approximately $2.34\%$ and $23.18\%$ of those consumed by DS-QAOA+ and QLS, respectively. The number of multi-qubit controlled gates consumed is $0.49\%$ and $7.36\%$ of those consumed by DS-QAOA+ and QLS, respectively.
  • On the 3-regular graphs, the number of iterations consumed by PQA is approximately $6.39\%$ and $65.92\%$ of those consumed by DS-QAOA+ and QLS, respectively. The amount of runtime consumed by PQA is approximately $6.43\%$ and $61.49\%$ of those consumed by DS-QAOA+ and QLS, respectively. The number of qubits consumed by PQA is approximately $2.17\%$ and $21.76\%$ of those consumed by DS-QAOA+ and QLS, respectively. The number of multi-qubit controlled gates consumed is $0.29\%$ and $2.88\%$ of those consumed by DS-QAOA+ and QLS, respectively.
V Conclusion and outlook
In this paper, we propose a heuristic quantum algorithm termed PQA to reduce the number of qubits in the application of QAOA+ when solving the MIS problem. The goal of PQA is to progressively construct a subgraph that is likely to contain the MIS solution of the original graph and then get the approximate solution by solving the MIS problem on this subgraph. To induce such a subgraph, PQA starts with a small-scale initial subgraph and progressively expands its graph size utilizing heuristic expansion rules. After each expansion, PQA solves the MIS problem on the current subgraph using QAOA+. In each run, PQA repeats the graph expansion and solving process until a predefined stopping condition is reached. The numerical results demonstrate that PQA can achieve higher OARs and AARs than DS-QAOA+ and QLS under the same optimization runs and level depths, indicating that PQA gets higher solution quality and has greater stability. Additionally, PQA is more efficient than DS-QAOA+ and QLS, which is reflected that PQA consumes fewer qubits and multi-qubit controlled $R_{x}$ gates to achieve the same AR as DS-QAOA+ and QLS. Given that NISQ devices are characterized by limited qubit connectivity and high gate noise, the reduction in qubit and multi-qubit controlled gate consumption in PQA significantly enhances its feasibility for near-term quantum hardware. In our work, the heuristic expansion strategies in PQA prioritize nodes that increase the likelihood of retaining the MIS solution while reducing additional quantum resource consumption. To maintain fine-grained control over subgraph growth, PQA starts with a small subgraph and incrementally expands it by adding one node at a time during each expansion step. However, this incremental approach may lead to a higher number of optimization rounds, potentially increasing the overall computational overheads. To address this limitation and improve convergence efficiency, future research should explore more effective graph construction strategies and the optimal choice of initial size, taking into account the graph structure to maintain PQA performance. One possible approach to reducing external optimization rounds is to introduce multiple vertices in each expansion step. However, further research is required to evaluate its impact on the effectiveness and stability of PQA. Another promising direction is to incorporate machine learning techniques to optimize the selection of initial graph size and dynamically refine the graph expansion strategy based on feedback obtained during the optimization process. These advancements could significantly reduce the number of iterations required in the external optimization loop, thereby enhancing the computational efficiency of PQA while improving its scalability to tackle larger problem instances. Finally, although this paper focuses on the MIS problem, the progressive solving framework can also be applied to other combinatorial optimization problems, such as the Maximum Weighted Independent Set problem, the Facility Location Problem, and the Knapsack Problem. A promising direction for future research is to design tailored expansion strategies for different optimization problems and investigate the applicability and performance of PQA in solving these problems. Acknowledgements Thanks for the Tianyan Quantum Computing Program. This work is supported by National Natural Science Foundation of China (Grant Nos. 62371069, 62372048, 62272056), and BUPT Excellent Ph.D. Students Foundation(CX2023123).
Appendix A The extension of the subgraph
Denote the current subgraph as $G_{i}=(V_{i},E_{i})$. For the first subgraph in FIG. A1, $V_{i}=\left\{0,1,2\right\}$ and $E_{i}=\left\{(0,1),(0,2)\right\}$. The candidate vertex set $V_{C}=V-V_{i}=\left\{3,4\right\}$. Our goal is to select the vertex from $V_{C}$ that has the least connection with the current subgraph to be added to $G_{i}$. Specifically, if node 3 is added to $G_{i}$, $C(v_{c}=3)=1$. Similarly, if node 4 is added, $C(v_{c}=4)=1$. Since both nodes have the same connection, we proceed to the next round of selection. When node 3 is added to $V_{i}$, the edge $(0,3)$ is included in $E_{i}$, leaving $V_{C}=\left\{4\right\}$ and $C(v_{c}=4)=2$. Similarly, if node 4 were added, $C(v_{c}=3)=2$. After evaluating both nodes, since they equally satisfy the selection criteria, one can be randomly added to $V_{i}$. If node 3 (or 4) is added, the edge $(0,3)$ (or $(0,4)$) is included in $E_{i}$. The resulting subgraph becomes is $G_{i+1}=\left\{\left\{0,1,2,3\right\},\left\{(0,1),(0,2),(0,3)\right\}\right\}$ (or $G_{i+1}=\left\{\left\{0,1,2,4\right\},\left\{(0,1),(0,2),(0,4)\right\}\right\}$).
Figure A1: The target graph and its induced subgraphs.
For the second subgraph in FIG. A1, $V_{i}=\left\{0,1,3\right\}$ and $E_{i}=\left\{(0,1),(0,3)\right\}$. The candidate vertex set $V_{C}=\left\{2,4\right\}$, with $C(v_{c}=2)=1$ and $C(v_{c}=4)=2$. We choose the node with the smallest $C(v_{c})$ value from the candidate vertex set. Thus, node 2 is the next node to be added to $V_{i}$, and the edge $(0,2)$ is added to $E_{i}$. The newly constructed subgraph is $G_{i+1}=\left\{\left\{0,1,2,3\right\},\left\{(0,1),(0,2),(0,3)\right\}\right\}$.
Appendix B The methods of calculating total resources
When the algorithm reaches a given AR at a certain level depth, the total resource consumption is calculated as the product of the total number of runs required to achieve the given AR under this level depth and the average resource consumption per run. Therefore, the key to this calculation lies in determining the required level depths, the total number of runs, and the resource consumption per run. To determine at which level depth the algorithm can achieve the given AR, we vary the level depth from 1 to $n$ and perform multiple optimization runs at each depth. Specifically, we conduct 100 optimization runs at each level depth $p$ for $p\leq 6$, and $2^{p}$ optimization runs when $p>6$. After completing all simulations, we first determine the shallowest level depths at which the algorithm achieves the given AR, based on the optimal approximate ratio obtained from multiple optimization runs at each depth. Next, we calculate the total number of runs and average resource consumption at this level depth. The total number of runs required to achieve the given AR is estimated using the following method. First, we calculate the number of runs $Q$ where the resulting AR exceeds the given threshold across multiple optimization runs. Next, we calculate the ratio of $Q$ to the total number of runs at level depth $p$, which represents the probability of exceeding the threshold. The reciprocal of this probability provides the total number of runs needed for the algorithm to achieve the given AR. We then calculate the total resource consumption for each algorithm to achieve the given AR on each graph. Next, we average the total number of runs, iterations, qubits, multi-qubit controlled $R_{x}$ gates, and runtime for each graph type, and use these values to represent the resources required by each algorithm to achieve the given AR on that graph type.
(a) Total runs, ER with prob = 0.4
(b) Total runs, ER with prob = 0.5
(c) Total runs, 3-regular graphs
Figure A2: The total optimization runs of various algorithms when they achieve the given AR on different graph types.
B.1 The methods of calculating the runtime.
@@ Appendix B The methods of calculating total resources
Runtime serves as a comprehensive measure of the execution efficiency of the algorithm, incorporating both quantum and classical computing time. For algorithms like QLS and PQA, classical computations (such as determining edge existence and graph expansion) contribute to the total runtime. In contrast, DS-QAOA+ excludes these computations. Although the time for a single-edge judgment is minimal, we include this classical computation time in our runtime comparison to maintain fairness. For QLS and PQA, the time required for a single optimization run includes both quantum and classical components. The classical time $T_{c}$ is the time consumed in determining whether there is an edge between vertices during updating the global solution or expanding the subgraph. Assuming that the time required for a single determination is $t_{j}$, and the total number of determinations required by the algorithm in an optimization run is $J$, then
$T_{c}=J\times t_{j}.$ (A1)
@@
The quantum running time $T_{q}$ includes the time required for initial state preparation, quantum state measurements, and quantum gate evolution across multiple iterations runtime . In each iteration, $M$ repeated measurements are needed to calculate the average expectation value of the objective function. The time consumed in a single repetition is
$t_{q}=t_{p}+t_{M}+(\text{circuit depth})\times t_{G},$ (A2)
@@
$T_{q}=T_{\text{ITR}}\times\text{ITRs},$ (A4)
where ITRs are the number of iterations required for the objective function to converge in this run. Therefore, the time consumed in a single run for QLS and PQA is
$T_{\text{PQA}}=T_{\text{q, PQA}}+T_{\text{c, PQA}},$ (A5)
$T_{\text{QLS}}=T_{\text{q, QLS}}+T_{\text{c, QLS}},$ (A6)